The salt cesium sulfate is soluble in water. When 20.9 g Cs2SO4 is dissolved in 112.00 g water, the temperature of the solution decreases from 25.00 °C to 22.81 °C. Based on this observation, calculate the dissolution enthalpy, ΔdissH, of Cs2SO4.

Assume that the specific heat capacity of the solution is 4.184 J g-1 °C-1 and that the energy transfer to the calorimeter is negligible.

ΔdissH = _____ kJ/mol
Answer #1

Moles of Cs2So4 = 20.9 g / 361.87 gmol-1 = 0.05775 mol

Heat of dissoloution = Heat absorbed by water

Heat released  by Water = m * C* delta T

mass of water m = 112 g

change in temperature = 25- 22.81 = 2.19 deg

c = 4.184 J / g °C

so heat released by water = 112 g * 2.19 °C * 4.184 J/ g °C = 1026.25 J

= 1026.25 J x 1 KJ / 1000 J = 1.026 KJ  

so heat of dissolution per mole = 1.026 KJ   / 0.05775 mol   = 17.8 Kj/ mol